OpenStudy (anonymous):

Integral Question: Find the average value of the function f on the given interval f(x) = (1/x^2) [1,5]

OpenStudy (anonymous):

are you just asking to integrate that function? I'm not sure what you mean by average value

OpenStudy (anonymous):

Average value is when you take the term and multiply it by 1/(b-a) or just f(c) = b-a

OpenStudy (anonymous):

Ah one of the calc1 theorems, sorry I can't remember those off the top of my head.

OpenStudy (anonymous):

\[\frac{ 1 }{ b-a }\int\limits_{a}^{b}f(x)dx\]

OpenStudy (anonymous):

So far I got: The integral for 5 to 1 (x^-2) = x^-1/1

OpenStudy (anonymous):

That is the average value of a function, using calculus

OpenStudy (anonymous):

Correct Andras. Did I integrate well so far? I don't want to get confused with the signs and negative exponents.

OpenStudy (anonymous):

This integral is a polynomial. So I really hope that you can do the integral

OpenStudy (anonymous):

You are missing a minus sign!

OpenStudy (anonymous):

Oh yeah lol my teacher keeps reinforcing that you can't take the area of a polynomial. I got -x^-1 = -1/5 - 1 = -1/5 -5/5 = -6/5

OpenStudy (anonymous):

Then would I just multiply it by 1/(b-a), which is 1/4 ? -6/5 multiplied by 1/4 = -6/20 = -3/10 That's what I got but I wanted to confirm it.

OpenStudy (anonymous):

Sorry had to leave for a sec, Im back now.

OpenStudy (anonymous):

f(x) = (1/x^2) Think!!!! Can this: -3/10 be the average value of the function between 1 and 5??

OpenStudy (anonymous):

Oh no it can't. It's not in the domain. Where did I go wrong?

OpenStudy (anonymous):

I got -x^-1 for the integral. You are correct But when you put the values in make sure you are treating all the - signs correctly. Will you give it another go?

OpenStudy (anonymous):

Yeah sure just half a minute.

OpenStudy (anonymous):

It should be -1/5 +1 right? Which is -1/5 + 5/5 = 4/5

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Oh wow oh wow. 4/5(1/4) = 4/20 = 1/5 Is that in the domain?

OpenStudy (anonymous):

Yeah that is it. Not the domain! The question should be: Is that in the range?

OpenStudy (anonymous):

Gotcha :) Thank you so much!